# Multi-Type Point Cloud Autoencoder: A Complete Equivariant Embedding for Molecule Conformation and Pose

### v1

@@ Abstract
The point cloud is a flexible representation for a wide variety of data types, and is a particularly natural fit for the 3D conformations of molecules. Extant molecule embedding/representation schemes typically focus on internal degrees of freedom, ignoring the global 3D orientation. For tasks that depend on knowledge of both molecular conformation and 3D orientation, such as the generation of molecular dimers, clusters, or condensed phases, we require a representation which is provably complete in the types and positions of atomic nuclei and roto-inversion equivariant with respect to the input point cloud. We develop, train, and evaluate a new type of autoencoder, molecular O(3) encoding net (Mo3ENet), for multi-type point clouds, for which we propose a new reconstruction loss, capitalizing on a Gaussian mixture representation of the input and output point clouds. Mo3ENet is end-to-end equivariant, meaning the learned representation can be manipulated on O(3), a practical bonus for downstream learning tasks. An appropriately trained Mo3ENet latent space comprises a universal embedding for scalar and vector molecule property prediction tasks, as well as other downstream tasks incorporating the 3D molecular pose.
@@ 1 Introduction
Point clouds are an extremely flexible data type used in many 3D modelling applications [1]. While often used as a primitive representation for the surface of some object [2], point clouds are also a natural representation for molecular information [3, 4, 5]. Indeed, in the most common file formats for 3D atomistic molecular data encode a simple list of points and associated atom types, that is, a multi-type point cloud. Despite extensive prior work in both point cloud modelling and molecular representations, no extant model explicitly decomposes 3D molecular structures to a fixed-size equivariant latent embedding, and then reconstructs them in atomistic detail without any scaffolding, sampling, or other external information.
@@
Such an autoencoder would be a useful tool in several problem areas. The perhaps most exciting application is in the growing field of machine learning for molecular materials design [6, 7, 8, 9, 10, 11, 12, 13, 14, 15], where one may need to condition a downstream task on the precise geometry and rotational orientation of an input molecule, in the form of an embedding vector prepared by the encoder.
@@
Further, such an encoder represents the ‘ultimate’ brute-force answer to the molecule representation learning problem. Currently, it is common practice to co-train molecule representation and property prediction models [16, 17] in the form of graph neural networks, although there is also impressive recent work on extensively pretrained representation models [18]. Co-training is generally an efficient approach, allowing one to make good use of finite datasets. However, in the limit of large datasets and a well-converged pretrained autoencoder model, we can obviate the need for complex and sometimes challenging-to-train graph neural networks, and train simple and reliable multilayer perceptrons directly on our encoder embedding.
Finally, the decoder half of this autoencoder would be a new and potentially interesting generative model for molecular data, with property-driven sampling and search enabled by conditional training and careful engineering of the embedding space. In the approach presented here, the decoder has the additional advantage of sampling in a single forward pass of the model, as opposed to modern diffusion approaches which may require a number of forward passes to generate a single sample [19, 20, 21, 22].
@@
With our work, we advance the field of point cloud learning by developing a new reconstruction loss and model architecture for multi-type point clouds which is smooth and fast to compute, returns trainable gradients, and allows for varying input composition and size. We leverage developments in equivariant graph learning for point-structured data [23, 4, 24], to develop and optimize an end-to-end equivariant autoencoder model. If accurately converged, the latent O(3) equivariant embedding between encoder and decoder in this model must necessarily contain the complete 3D coordinate and type information about a given molecule, and therefore serve as a universal embedding space for molecular data. Such an embedding could be leveraged for any downstream task dependent on molecular structure and/or pose, without fear of common issues in graph learning such as lack of representational power [3], with the added benefit of rotatability.
@@
Our Mo3ENet model comprises an equivariant graph neural network (EGNN) encoder, which embeds an input point cloud as a fixed dimensional equivariant representation $\vec{g}_{k\times 3}$, and an equivariant multilayer perceptron (EMLP) decoder. The encoder is flexible in the number of input points, and the output returns a fixed-size ‘swarm’ of output points. Using a ‘swarm’ decoder with a fixed number of output points enables us to use a simple and efficient MLP-type model as the decoder, rather than a more complex graph or sequence model. A large swarm size with learnable weights ensures flexibility of the decoding with respect to input size and composition. Both inputs and outputs are mapped to Gaussian mixtures (GMs), and their pairwise overlaps computed to determine the reconstruction fidelity. Training is facilitated by an reduction of the GMs’ widths until neighboring points no longer significantly overlap, and the distance between input and output distributions is minimized, corresponding to perfect reconstruction of the input point cloud.
As an illustration of our autoencoder model in a chemistry context, we use a large and diverse set of molecules as a source of training data and assess its representation ability on this dataset on equivariant reconstruction and property prediction tasks.
@@ 2 Related work
Given their utility, there is vast prior literature on point cloud learning, including object classification, semantic segmentation, generative modelling, representation learning, and much more [25, 26, 2, 27, 28, 29]. Various model architectures have been proposed to process this data type, with inputs cast as voxel grids, meshes, and real-space graphs. Models incorporating underlying physical equivariances into their building blocks have also seen increased uptake in recent years [30, 24, 31, 32], with particular popularity in the chemical modelling community [33, 23, 4]. In this vein, several generative approaches including various autoencoders [2, 34] and generative adversarial networks [35, 36], have been developed, with particular focus on point cloud reconstruction. While many insights from this field, and particularly the original PointNet architecture [37], have been influential in molecular modelling, there are as yet limited tools for the analysis and generation of point clouds with multiple types.
@@
The study of molecule representations goes back decades in the chemistry literature. Such approaches run the gamut from straightforward and interpretable, such as the bag of bonds [38] or SMILES strings, to the more structurally-informed Coulomb matrix [39], to physics-inspired descriptors [40, 41], and on to high-dimensional learned embeddings [42, 18, 43, 44].
The most powerful pretrained approach currently appears to be the Uni-Mol framework [18], which combines a large and powerful graph model, self-supervised learning, property prediction, and very large datasets, to train a high-quality molecule representation. Despite impressive benchmark performance, however, Uni-Mol does not explicitly encode and decode the complete molecular 3D structure.
@@ 3 Models and methods
3.1 Training
We trained on the QM9 dataset [52, 53], containing 133k molecules with up to 9 heavy atoms, including H, C, N, O, and F. We optimized models including and excluding hydrogen atoms, both to compare performance and because in many relevant datasets such as the Cambridge Structural Database [54] of molecular crystals, hydrogen positions are inconsistent or absent. Insights from training on QM9 should be transferable to training on larger datasets with larger molecules, e.g., GEOM [55], QMugs [56]. The dataset was split 80:20 into train:test to assess generalization to unseen molecules.
We also attempted training with point clouds from a random uniform distribution in both coordinates and atom types, but found negligible benefit to model performance on molecular data. This should, perhaps, not be surprising, since molecular information is structured by certain common correlations, such as bond lengths and angles, which the autoencoder can exploit to more easily reconstruct the input. Random point clouds share none of these correlations and, thus, are far outside the distribution of molecular data.
@@ 3.2 Model architecture
Figure 1: (a) Overall summary of autoencoder architecture. Molecule point cloud coordinates, $\vec{r}$ and atom types $z$ are embedded and bottlenecked to dimension $f_{emb}$ by the encoder. Then, the decoder transforms the embedding to $n_{j}$ independently weighted points in the $3+1+n_{c}$ dimensional output cloud, where $n_{c}$ is the number of classes or types in the dataset and one extra dimension for the learned weights. (b) Outline of the architecture of the equivariant graph model which comprises the encoder, outputting the equivariant embedding $\bm{\vec{g}}$ and its scalarized counterpart $\bm{\tilde{g}}$.
@@
The autoencoder is comprised of an encoder and decoder, inputting and outputting scalars $\bm{s}$ and vectors $\bm{\vec{v}}$. In the molecular context, the initial nodewise scalar inputs $\bm{s}$ are the atom types and radial distances from the molecule centroid, and vector inputs $\bm{\vec{v}}$ are the Cartesian point vectors from the molecule centroid. The model is end-to-end equivariant on O(3), that is, rotations and inversions of the input result in identical rotations to the embedding and decoding. As illustrated in Figure 1(a), during a forward pass, the atoms in a given molecule are embedded as nodes in a graph, nonlinearly transformed, and aggregated to a fixed size latent embedding, $\bm{\vec{g}}$, which is decoded to a swarm of points $\bm{\vec{d}}$.
@@
The primary building blocks of our model are an equivariant multilayer perceptron (EMLP), graph convolution layers, and graph-wise aggregators. In constructing these modules, we borrow concepts from prior works on equivariant models on scalars and vectors [23, 4, 24], which process directly the point positions in Cartesian space. Only a few transformations of vector features retain O(3) equivariance; we nevertheless realize a flexible model using the following basic operations: (1) linear combination, $\vec{v}^{t+1}=\sum_{j}c_{j}\vec{v}^{t}_{j}$, (2) adjustment of vector lengths via multiplication by scalars, and (3) vector embedding and activation, similar to the projection scheme developed in [24], as well as the usual neural network operations for scalars. Full technical details of model construction are provided in the supplemental material (SM).
@@
The encoder model, shown in Figure 1(b) is comprised of an embedding step, alternating EMLP and graph convolution modules, and finally graph aggregation. A ‘scalarizing’ module $\mathcal{S}$, can also be used to generate a rotation invariant scalar representation of the full vector embedding. The decoder is a straightforward EMLP model, returning scalar and vector features of the output swarm.
$\begin{split}\bm{s}^{N},\vec{\bm{v}}^{N}&=\mathrm{EMLP}(\tilde{\bm{g}},\vec{\bm{g}})\\ \vec{\bm{d}}&=\vec{\bm{v}}^{N}\|\bm{s}^{N}\end{split}$ (1)
@@
$\vec{\bm{d}}$ is a vector with shape $n_{j}\times(3+n_{c}+1)$, corresponding to the $n_{j}$ output swarm nodes, the 3 Cartesian dimensions, the $n_{c}$ classwise probabilities, and the raw amplitude of the nodewise weight. The decoder is equivariant with respect to the Cartesian dimensions and invariant with respect to type and weight. The $n_{c}$ elements are normed via softmax within each node to a particle-type probability. Each output point also outputs a learnable weight $w_{j}$ computed as the softmax over the predicted amplitude for all points in the output swarm, multiplied by the number of atoms in the input graph, such that $\sum_{j}w_{j}=n_{i}$.
@@ 3.3 Reconstruction loss
To train a multi-type point cloud autoencoder, we develop a reconstruction loss function with the following properties: smoothly converging in particle positions and type variations, fast to evaluate, permutation invariant, scales easily to large numbers of particles and particle types, and provides a trainable gradient.
@@
In the convergent limit, diagrammed in Figure 2, each point in the input should be exactly overlapped by the GM of the output. The input GM is centred on the Cartesian points in the input cloud and one-hot encoded in the $n_{c}$ type dimensions, with a variable number of points $n_{i}$. The output GM is likewise constructed, centred in the points in the output distribution, with learned probabilities for each particle type dimension, a fixed number of points $n_{j}$ predicted by the decoder, and appropriately normed to a total probability mass matching the input GM. The output comprises a fixed number of points in a ‘swarm’, $n_{j}$, with $n_{j}$ in practice significantly greater than $n_{i}$.
The decoder could be designed as a graph model, mirroring the encoder, though we found empirically that this type of model failed to train, and required auxiliary information outside the fixed-size latent embedding (i.e., the number of atoms in the molecule). Instead, considering the decoder as an EMLP simplifies the architecture and demonstrates efficient learning, afforded by the fact that our GM reconstruction loss does not require the number of points in input and output distributions to be equal.
The overlap between points $i$ and $j$ in the input and output, respectively, is simply given by their Gaussian distributed euclidean distances,
$A_{ij}(\vec{x}_{i},\vec{y}_{j})=w_{j}\cdot e^{\frac{-d_{ij}^{2}}{\sigma}},$ (2)
with $d_{ij}$ as the euclidean distance between points $\vec{x}_{i}$ and $\vec{y}_{j}$, $\vec{x}_{i}=\vec{r}_{i}\|e_{i}$, $\vec{r}_{i}$ the 3D coordinate, and $e_{i}=c_{T}\cdot\mathrm{OneHot}(z_{i})$ a one-hot encoding for the type of particle $i$. $w_{j}$ is, again, the learned weight for output node $j$. The predicted points $j$ are similarly encoded as $\vec{y}_{j}=\vec{r}_{j}\|\left(c_{T}\cdot P(z_{j})\right)$, with the normalized probabilities over particle types substituted for the one-hot encoding. $c_{T}$ is a scaling factor which adjusts the relative distances between particle types. It is a tunable hyperaparamter which can have significant impact on the training dynamics if taken too large or too small compared to $\sigma$. In principle, we expect to see good performance when $c_{T}$ is taken on the order of interatomic distances, typically 2Å, such that atoms of different types are neither too close, such that they overlap, or too far, such that the type space becomes too sparse.
@@
In practice, we initialize $\sigma$ somewhat large to ensure the decoder receives nonzero losses and therefore gradients at the beginning of training. In such early stages, the loss is the difference between the input-output overlap and the input self-overlap, from adjacent points’ Gaussians intruding on one-another,
@@
A converging series of decoder output GMs for a given molecule is shown in Figure 3. As the reconstruction loss improves, $\sigma$ is reduced, increasing the problem difficulty, until the overlap between points in the input becomes negligible.
We experimented with several alternate or auxiliary losses, for example, one enforcing the correct distribution of atom types in the decoder output, with the goal of accelerating model convergence. In practice, however, these did not generally improve model performance, possibly due to limiting the optimization pathways for the basic reconstruction loss. Two exceptions were employed to ensure robust gradients in early training.
$L_{c}=\sum_{j}^{n_{j}}\mathrm{Max}\{(\lVert\vec{r}_{j}\rVert-1),0\}$ (5)
constrains output points to be within the unit sphere of the origin, which is the maximum span of all the input point clouds. This prevents decoded points straying too far from the range of the input, which would lead to vanishingly small gradients from the reconstruction loss. Secondly, since we empirically observed some point weights vanishing in early epochs and inhibiting convergence, we implement a constraint on the minimum allowable output point weight $w_{j}$, at 1% of the mean for a given output swarm. In practice, both constraining losses converge trivially and do not seem to constrain optimization against the reconstruction loss.

### v3

@@ Abstract
Representations are a foundational component of any modelling protocol, including on molecules and molecular solids. For tasks that depend on knowledge of both molecular conformation and 3D orientation, such as the modelling of molecular dimers, clusters, or condensed phases, we desire a rotatable representation that is provably complete in the types and positions of atomic nuclei and roto-inversion equivariant with respect to the input point cloud. In this paper, we develop, train, and evaluate a new type of autoencoder, molecular O(3) encoding net (Mo3ENet), for multi-type point clouds, for which we propose a new reconstruction loss, capitalizing on a Gaussian mixture representation of the input and output point clouds. Mo3ENet is end-to-end equivariant, meaning the learned representation can be manipulated on O(3), a practical bonus. An appropriately trained Mo3ENet latent space comprises a universal embedding for scalar and vector molecule property prediction tasks, as well as other downstream tasks incorporating the 3D molecular pose, and we demonstrate its fitness on several such tasks.
@@ 1 Introduction
The question of how to represent a molecule to a learning model is the crucial first step in any modelling routine. A representation must be sufficiently rich to capture the relevant features, yet not too complex, which leads to overfitting. 3D molecular structures are typically cast as point clouds, i.e., lists of coordinates and atom types, which, by necessity, have varying sizes. Despite extensive prior work in both point cloud modelling and molecular representations, no extant model explicitly decomposes 3D molecular structures to a fixed-dimensional equivariant embedding (not dependent on the input graph size), and then reconstructs them in atomistic detail without any scaffolding, sampling, or other external information, as would, for example, an autoencoder model.
@@
Such an autoencoder would be a useful tool in several application areas. The perhaps most exciting of these is in the growing field of machine learning for molecular materials design [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], where one may need to condition a downstream task on the precise geometry and rotational orientation of an input molecule, represented by an embedding vector prepared by the encoder. Further, such an encoder represents a complete, if brute-force, solution to the molecule representation learning problem.
@@
Currently, it is common practice to co-train molecule representation and property prediction models [11, 12] with graph neural networks, although there is also impressive recent work on extensively pretrained representation models [13, 14, 15, 16, 17, 18]. Co-training is generally an efficient approach, allowing one to make good use of finite datasets. However, in the limit of large datasets and a well-converged encoder model, we can obviate the need for graph prediction models, with their attending computational and theoretical complexities [19, 20], and train simpler models such as multilayer perceptrons directly on a pretrained embedding.
@@
In this work, we advance the fields of point cloud learning and molecule representation by developing a new reconstruction loss and model architecture for multi-type point clouds which is smooth and fast to compute, returns trainable gradients, and allows for varying input composition and size. We leverage developments in equivariant graph learning for point-structured data [21, 22, 23] to develop and optimize an end-to-end equivariant autoencoder model. If well-converged, the latent O(3) equivariant embedding between encoder and decoder in this model must necessarily contain the complete 3D coordinate and type information about a given molecule, and therefore serve as a universal embedding space for molecular data. Such an embedding could be leveraged for any downstream task dependent on molecular structure and/or pose, without fear of common issues in graph learning such as lack of representational power [24], with the added benefit of rotatability.
Although we do not explore the possibility here, the decoder half of this autoencoder could be developed into a new generative model (variational autoencoder) for atomistic data. Property-driven sampling and search would then be possible via by conditional training and careful engineering of the embedding space. In the approach presented here, the decoder has the additional advantage of sampling in a single forward pass of the model, as opposed to modern diffusion approaches which may require a number of forward passes to generate a sample [25, 26, 27, 28].
@@
Our Mo3ENet model comprises an equivariant graph neural network (EGNN) encoder, which embeds an input point cloud as a fixed dimensional equivariant representation $\vec{g}_{k\times 3}$, with $k$ the embedding dimension, and an equivariant decoder, structured either as an EGNN or equivariant multilayer perceptron (EMLP). The encoder is flexible in the number of input points, and the output returns a fixed-size ‘swarm’ of output points. A large swarm size with learnable member weights ensures flexibility of the decoding with respect to input size and composition. Both inputs and outputs are mapped to Gaussian mixtures (GMs), with their pairwise overlaps computed to determine the reconstruction fidelity. Training is facilitated by an reduction of the GMs’ widths until neighboring atoms no longer significantly overlap, and the distance between input and output distributions is minimized, corresponding to perfect reconstruction of the input point cloud.
To demonstrate our autoencoder model in a chemistry context, we train it on a very broad distribution of randomly generated molecular conformers, based on QM9 and ZINC22 [29, 30, 31] molecular backbones, and assess its representation ability on molecule reconstruction and property prediction tasks, as well as demonstrate that, combined with unit cell parameters, the molecule embedding can be used to predict crystal properties such as lattice energies. This last task shows the plausibility of more complex modelling tasks, including molecular crystal search and optimization.
@@ 2 Related work
Given their utility, there is vast prior literature on point cloud learning, including object classification, semantic segmentation, generative modelling, representation learning, and much more [32, 33, 34, 35, 36, 37]. Various model architectures have been proposed to process this data type, with inputs cast as voxel grids, meshes, and real-space graphs. Models incorporating underlying physical equivariances into their building blocks have also seen increased uptake in recent years [38, 23, 39, 40, 41, 42], with particular popularity in the chemical modelling community [43, 21, 22]. In this vein, several generative approaches including various autoencoders [34, 44] and generative adversarial networks [45, 46], have been developed, with particular focus on point cloud reconstruction. While many insights from this field, and particularly the original PointNet architecture [47], have been influential in molecular modelling, there are as yet limited tools for the analysis and generation of point clouds with multiple types.
@@
The study of molecule representations goes back decades in the chemistry literature. Such approaches run the gamut from straightforward and interpretable, such as the bag of bonds [48] or SMILES strings, to the more structurally-informed Coulomb matrix [49], to physics-inspired descriptors [50, 51], and on to high-dimensional learned embeddings [52, 13, 53, 54, 16, 15, 14].
The most powerful pretrained approach currently appears to be the Uni-Mol framework [13], which combines a large and powerful graph model, self-supervised learning, property prediction, and very large datasets, to train a high-quality molecule representation. Despite impressive benchmark performance, however, Uni-Mol does not explicitly encode and decode the full molecular point cloud, instead, predicting positions and types of single corrupted or masked atoms.
@@ 3.1 Model architecture
Figure 1: (a) Overall summary of autoencoder architecture. Molecule point cloud coordinates, $\vec{r}$ and atom types $z$ are embedded and bottlenecked to dimension $f_{emb}$ by the encoder. Then, the decoder (an equivariant MLP or GNN with fixed graph size) transforms the embedding to $n_{j}$ independently weighted points in the $13+n_{c}+$ 1 dimensional output cloud, where $n_{c}$ is the number of classes or types in the dataset and one extra dimension for the learned weights. (b) Outline of the architecture of the encoder graph model, outputting the equivariant embedding $\bm{\vec{g}}$ and its scalarized counterpart $\bm{\tilde{g}}$.
@@
The Mo3ENet autoencoder is comprised of an encoder and decoder, each inputting and outputting scalars $\bm{s}$ and vectors $\bm{\vec{v}}$. In the molecular context, the initial nodewise scalar inputs $\bm{s}$ are the atom types and radial distances from the molecule centroid, and vector inputs $\bm{\vec{v}}$ are the Cartesian point vectors from the molecule centroid, together comprising a single molecular graph. The model is end-to-end equivariant on O(3), that is, rotations and inversions of the input result in identical rotations to the embedding and decoding. As illustrated in Figure 1(a), during a forward pass, the atoms in a given molecule are embedded as nodes in a graph, nonlinearly transformed, and aggregated to a fixed size latent embedding, $\bm{\vec{g}}$, which is decoded to a swarm of points $\bm{\vec{d}}$.
@@
The primary building blocks of our model are EMLPs, graph convolution layers, and graph-wise aggregators. In constructing these modules, we borrow concepts from prior works on equivariant models on scalars and vectors [21, 22, 23], which process directly the point positions in Cartesian space. Only a few transformations of vector features retain O(3) equivariance; we nevertheless realize a flexible model using the following basic operations: (1) weighted linear combination, (2) adjustment of vector lengths via multiplication by scalars, and (3) vector embedding and activation, similar to the projection scheme developed in [23], as well as the usual neural network operations for scalars. Full technical details of model construction are provided in the appendix.
@@
The encoder model, shown in Figure 1(b) is comprised of an embedding step, alternating EMLP and graph convolution modules, and finally, graph aggregation. A ‘scalarizing’ module $\mathcal{S}$, can also be used to generate a rich, rotation invariant scalar representation of the full vector embedding. The decoder can be either another equivariant graph model (sans global aggregation), or an equivariant MLP, returning scalar and vector features of the output swarm,
$\begin{split}\bm{s}^{N},\vec{\bm{v}}^{N}&=\mathrm{Decoder}(\tilde{\bm{g}},\vec{\bm{g}})\\ \vec{\bm{d}}&=\vec{\bm{v}}^{N}\|\bm{s}^{N},\end{split}$ (1)
with the EGNN decoder first generating scalar and vector node embeddings, and positions from the encoder embedding, for a constant number of nodes, via a learned linear transformation. $\vec{\bm{d}}$ is a vector with shape $n_{j}\times(3+n_{c}+1)$, corresponding to the $n_{j}$ output swarm nodes, the 3 Cartesian dimensions, the $n_{c}$ atom type probabilities, and the raw amplitude of the nodewise weight.
@@
The decoder is equivariant with respect to the Cartesian dimensions and invariant with respect to type and weight. The $n_{c}$ elements are normed via softmax within each node to a particle-type probability. Each output point also includes a learnable weight $w_{j}$ computed as the softmax over the predicted amplitude for all points in the output swarm for a given molecule, multiplied by the number of atoms in the input graph, such that $\sum_{j}w_{j}=n_{i}$. Note that throughout this work, $n_{i}$ corresponds to the number of nodes in the input molecular graph, and $n_{j}$ the number of nodes in the decoded output swarm.
Given a constant graph size in the decoder, the choice of EMLP or EGNN is not obviously theoretically indicated, and indeed, a graph model introduces extra complexity. However, empirically, the graph version is observed to have superior convergence properties, and so the results in Section 2 are generated using the EGNN version. All model hyperparameters and further training details are given in the appendix.
@@ 3.2 Reconstruction loss
In order to train a multi-type point cloud autoencoder, we develop a reconstruction loss function with the following properties: smoothly converging in particle positions and type variations, fast to evaluate, permutation invariant, scales easily to large numbers of particles and particle types, and provides a trainable gradient.
@@
In the convergent limit, diagrammed in Figure 2, each point in the input should be exactly overlapped by the GM of the output. The input GM is centred on the Cartesian points in the input cloud and one-hot encoded in the $n_{c}$ type dimensions, with a variable number of input points $n_{i}$. The output GM is constructed likewise, centred in the points in the output distribution, with learned probabilities for each particle type dimension. The output comprises a fixed number of points $n_{j}$, with $n_{j}$ always greater than $n_{i}$.
The overlap between points $i$ and $j$ in the input and output, respectively, is given by their Gaussian distributed distances,
$A_{ij}(\vec{x}_{i},\vec{y}_{j})=w_{j}\cdot e^{-\frac{d_{ij}^{2}}{\sigma}},$ (2)
with $d_{ij}$ as the euclidean distance between points $\vec{x}_{i}$ and $\vec{y}_{j}$, $\vec{x}_{i}=\vec{r}_{i}\|e_{i}$, $\vec{r}_{i}$ the 3D coordinate, and $e_{i}=c_{T}\cdot\mathrm{OneHot}(z_{i})$ a one-hot encoding for the type of particle $i$. $w_{j}$ is, again, the learned weight for output point $j$. The predicted points $j$ are similarly encoded as $\vec{y}_{j}=\vec{r}_{j}\|\left(c_{T}\cdot P(z_{j})\right)$, with the normalized probabilities over particle types substituted for the one-hot encoding. $c_{T}$ is a scaling factor that adjusts the relative distances between particle types. It is a tunable hyperaparamter that can have significant impact on the training dynamics if taken too large or too small compared to $\sigma$. In principle, we expect to see good performance when $c_{T}$ is taken on the order of interatomic distances, typically 2Å, such that atoms of different types are neither too close, such that they overlap, or too far, such that the type space becomes too sparse.
@@
This functional form attracts probability mass towards each input particle when its overlap with the outputs is less than 1, and repels it when the total overlap is greater than 1.
In practice, we initialize $\sigma$ somewhat large to ensure all swarm members receive finite gradients at the beginning of training. In such early stages, the loss is the difference between the input-output overlap and the input self-overlap, from adjacent points’ Gaussians intruding on one-another,
@@
A converging series of decoder output GMs for a given molecule is shown in Figure 3. As the reconstruction loss improves, $\sigma$ is reduced, increasing the problem difficulty, until the overlap between points in the input becomes negligible. The $\sigma$ reduction schedule, and specifically the threshold loss at which we trigger a reduction, is a critical hyperparameter affecting the rate of model convergence. If this threshold is set low, the model is required to have nearly perfect overlap at a given $\sigma$ before narrowing, meaning the relevant Gaussian components are very close to the atoms. Since the gradients at the peak of a Gaussian vanish, convergence in this regime can be extremely slow. Empirically, a reconstruction loss threshold of 0.15 was found to optimize the speed of convergence, supplying sufficient gradients without stranding output GM components too far from input atoms when $\sigma$ is narrowed.
We experimented with several alternate or auxiliary losses, for example, one enforcing the correct distribution of atom types in the decoder output, with the goal of accelerating model convergence. In practice, however, these did not generally improve model performance, possibly due to limiting the optimization pathways for the basic reconstruction loss. Three exceptions were employed to ensure robust gradients and no ‘orphaned’ nodes.
$L_{c}=\sum_{j}^{n_{j}}\mathrm{Max}\{(\lVert\vec{r}_{j}\rVert-r_{mol}),0\}$ (5)
constrains output points to be within the maximum radius of a given molecule. This prevents decoded points straying too far from the range of the input, which would lead to vanishingly small gradients from the reconstruction loss. Secondly, since we observed some point weights vanishing in early epochs and inhibiting convergence, we implement a penalty on positive deviations of the output point weights $w_{j}$ from the mean for a given output swarm. In practice, both constraining losses converge trivially and do not seem to haper optimization against the reconstruction loss. Finally, we impose a penalty on the distance of atom $i$ to the nearest swarm member $j$, which when applied gently helps keep atoms from getting missed entirely by the swarm, say if they are too far away for the gaussian overlap loss to develop a finite gradient.
$L_{d}=c_{d}\sum_{i}^{n_{i}}\min_{j}d_{ij}^{2}.$ (6)
where $c_{d}$ is a hyperparameter. At large values of $c_{d}$, training is significantly slowed, so we take in practice $c_{d}=0.1$.
The output dimension $n_{j}\times(4+n_{c})$ of the decoder model is typically high. Further, the loss calculation requires an all-to-all distance calculation between atoms and swarm nodes, naively scaling quadratically in both. We have trained EMLP decoders on molecules with up to 1024 swarm members, but have not noticed bottlenecking in training from these parts of the model. Perhaps for very large point clouds with many atom types, one might require more clever engineering to address these issues, but at this time this does not appear to be necessary.